OpenStudy (anonymous):

If T1(x) and T2(x) are onto linear transformations from Rn to Rm, then so is W(x) = T1(x) + T2(x). True or False? can you explain why?

OpenStudy (blockcolder):

Test for closure under addition and scalar multiplication.

OpenStudy (anonymous):

So T1(x) + T2(x) = T(x1+x2)? and cT1(x) +cT2(x) = c(T1(x)+T2(x))?

OpenStudy (anonymous):

If we know they are linear transformatoins, what's the point of checking for that?

OpenStudy (anonymous):

or do I just check only for W(x)?

OpenStudy (blockcolder):

No, you test W for closure. You check if it is true that W(x+y)=W(x)+W(y) and W(cx)=cW(x) by using the fact that T1 and T2 are linear transformations.

OpenStudy (anonymous):

But there is no W(x+y) for this problem. It's only W(x). So what do I add W(x) to?

OpenStudy (blockcolder):

If it's confusing, then you can just use different letters and instead prove that W(u+v)=W(u)+W(v) and W(cu)=cW(u), where u and v are vectors in R^n.

OpenStudy (anonymous):

OH, so it doesn't matter what vectors you put in there? You can have W(y) instead of W(x)?

OpenStudy (blockcolder):

Yeah.

OpenStudy (anonymous):

So say I did W(x+y), does that mean I would have T1(x)+T2(x) + T1(y)+T2(y)?